# LaFiCMIL: Rethinking Large File Classification from the Perspective of Correlated Multiple Instance Learning

### v3

@@ 4.2 LaFiCMIL
According to Theorem 1, we leverage Multi-layer Perceptron Rumelhart et al. (1986), BERT Devlin et al. (2018), LaFiAttention Layer and Layer Normalization Ba et al. (2016) as sub-functions to approximate the c-MIL score function $S$ defined in Equation 10.
Given a set of bags $\{X_{1},...,X_{N}\}$, where each bag $X_{i}$ contains multiple instances $\{x_{i,1},...,x_{i,n}\}$, a bag label $Y_{i}$, and a randomly initialized category vector $x_{i,category}$. The goal is to learn the maps: $\mathbb{X}\rightarrow\mathbb{T}\rightarrow\gamma$, where $\mathbb{X}$ is the bag space, $\mathbb{T}$ is the transformer space and $\gamma$ is the label space. The map of $\mathbb{X}\rightarrow\mathbb{T}$ can be defined as:
$\begin{split}X_{i}^{0}=&[x_{i,category};f(x_{i,1});...;f(x_{i,n})]\\ &+E_{pos},\indent X_{i}^{0},E_{pos}\in\mathbb{R}^{(n+1)\times d}\end{split}$ (11)
$\begin{split}Q^{l}=X_{i}^{l-1}W_{Q},&\indent K^{l}=X_{i}^{l-1}W_{K},\indent V^{l}=X_{i}^{l-1}W_{V},\\ &\indent l=1,...,L\end{split}$ (12)
where function $f$ is approximated by a pre-trained BERT model, $E_{pos}$ is the Positional Embedding , and $L$ is the number of $MSA$ block.
$\begin{split}head&=LaFiSA(Q^{l},K^{l},V^{l})\\ &=softmax(\frac{Q^{l}(\widetilde{K}^{l})^{T}}{\sqrt{d_{q}}})Z^{*l}softmax(\frac{\widetilde{Q}^{l}(K^{l})^{T}}{\sqrt{d_{q}}})V^{l},\\ \end{split}$ (13)
$MSA(Q^{l},K^{l},V^{l})=Concat(head_{1},...,head_{h})W_{O},$ (14)
$X_{i}^{l}=MSA(LN(X_{i}^{l-1}))+X_{i}^{l-1},\indent l=1,...,L$ (15)
where $W_{O}\in\mathbb{R}^{hd_{v}\times d}$, $head\in\mathbb{R}^{(n+1)\times d_{v}}$, $LaFiSA$ denotes the approximated Self-attention layer by Nyström method Baker (1977) according to Equation 8, $h$ is the number of head in each $MSA$ block, and Layer Normalization(LN) is applied before each $MSA$ block.
@@
0:  A set of bags (i.e., long documents) $\{X_{1},...,X_{N}\}$, a feature extraction function $f$ (i.e., BERT), a randomly initialized learnable category vector $x_{i,category}$.
0:
1:  for $X_{i}$ in $\{X_{1},...,X_{N}\}$ do
2:     divide the $i^{th}$ bag into instances: $\{x_{i,1};...;x_{i,n}\}\leftarrow X_{i}$
3:     $X_{i}^{0}\leftarrow[x_{i,category};f(x_{i,1});...;f(x_{i,n})]$
4:     $E_{pos}\leftarrow Positional\_Embedding(X_{i}^{o})$
5:     $X_{i}^{o}\leftarrow X_{i}^{o}+E_{pos}$
6:     parallelly computing MSA blocks: $l=1,...,L$
7:      $Q^{l}\leftarrow X_{i}^{l-1}W_{Q},K^{l}\leftarrow X_{i}^{l-1}W_{K},V^{l}\leftarrow X_{i}^{l-1}W_{V}$
8:      Compute landmarks from input $Q^{l}$ and landmarks
9:      from input $K^{l},\widetilde{Q}^{l}$ and $\widetilde{K}^{l}$ as the matrix form;
10:      Compute $\widetilde{F}\leftarrow softmax(\frac{Q^{l}(\widetilde{K}^{l})^{T}}{\sqrt{d_{q}}})$;
11:      Compute $\widetilde{B}\leftarrow softmax(\frac{\widetilde{Q}^{l}(K^{l})^{T}}{\sqrt{d_{q}}})$ ;
12:      Compute $\widetilde{A}\leftarrow softmax(\frac{\widetilde{Q}^{l}(\widetilde{K}^{l})^{T}}{\sqrt{d_{q}}})+;$
13:      $\hat{S}\leftarrow\widetilde{F}\times\widetilde{A}\times\widetilde{B}$
14:      $head\leftarrow\hat{S}V^{l}$
15:      $MSA^{l}\leftarrow Concat(head_{1},...,head_{h})W^{o}$
16:     for l in $\{1,...,L\}$ do
17:        $X_{i}^{l}\leftarrow MSA(LN(X_{i}^{l-1}))+X_{i}^{l-1}$
18:     end for
19:     $\hat{Y}_{i}\leftarrow MLP(LN((X_{i}^{L})^{(0)})),$
20:  end for
21:  Final predictions for documents $\{X_{1},...,X_{N}\}$ are $\{\hat{Y}_{1},...,\hat{Y}_{N}\}$
Algorithm 1 LaFiCMIL processing flow
@@
20NewsGroups multi-class 18 8461884618\,84618 846 10 1821018210\,18210 182 1132 7532 20 368.83$\pm$783.84 14.71
Book Summary multi-label 12 7881278812\,78812 788 10 2301023010\,23010 230 1279 1279 227 574.31$\pm$659.56 38.46
@@
EURLEX-57K multi-label 57 0005700057\,00057 000 45 0004500045\,00045 000 $6000$ $6000$ 4271 707.99$\pm$538.69 51.3
-Inverted multi-label 57 0005700057\,00057 000 45 0004500045\,00045 000 6000 6000 4271 707.99$\pm$538.69 51.3
Devign binary 27 3182731827\,31827 318 21 8542185421\,85421 854 $2732$ $7322$ 615.46$\pm$41 917.5441917.5441\,917.5441 917.54 39.76
From the above formulation, we can find that the most important part is to efficiently learn the map from bag space $\mathbb{X}$ to Transformer space $\mathbb{T}$. As illustrated in Figure 1, this map is approximated by a series of sub-functions which are approximated by various neural layers. The overall process is presented in Algorithm 1 and can be summarized as follows: given a large file, we use a BERT model to generate the representations of the divided chunks (i.e., instances in the concept of c-MIL). Then, we initialize a learnable category vector that follows a normal distribution and has the same shape as each instance. By considering the category vector as an additional instance, we learn the correlation between each instance using LaFiAttention layer. With the help of the attention mechanism, the category vector exchanges information with each chunk and extracts necessary features for large file classification. Finally, the category vector is fed into a fully connected layer to finalize the classification task.
5 Experiments
5.1 Experimental Setup
Datasets.
To ensure a fair comparison with baselines, we adopt the same benchmark datasets utilized in the state-of-the-arts for long document classification Park et al. (2022). We first evaluate LaFiCMIL on these six benchmark datasets: ① Hyperpartisan Kiesel et al. (2019), a compact dataset encompassing 645 documents, designed for binary classification. ② 20NewsGroups Lang (1995), comprising 20 balanced categories and 11 8461184611\,84611 846 documents.③ CMU Book Summary Bamman and Smith (2013), tailored for multi-label classification, contains 12 7881278812\,78812 788 documents and 227 genre labels. ④ Paired Book Summary Park et al. (2022), formulated by combining pairs of documents from the CMU Book Summary dataset, features longer documents. ⑤ EURLEX-57K Chalkidis et al. (2019), a substantial multi-label classification dataset consisting of 57 0005700057\,00057 000 EU legal documents and $4271$ available labels. ⑥ Inverted EURLEX-57K Park et al. (2022), a modified version of EURLEX-57K dataset in which the order of sections is inverted, ensuring that core information appears towards the end of the document. To better assess our method’s capability on handling extremely longer sequences, we also include a C programming language dataset Devign Zhou et al. (2019) for code defect detection, in which the long documents have extremely more tokens than the other six datasets.Table 1 offers detailed insights into the datasets, including metrics like average, maximum, and minimum token counts, along with the percentage of large documents, among other aspects. For a detailed description of each dataset, please refer to the comprehensive explanations in the Appendix A.1.
Implementation Details.
We split a long text document into chunks (i.e., c-MIL instances), and follow the standard BERT input length (i.e., 512 tokens) for each chunk. To ensure a fair comparison, in line with the baseline models in Park et al. (2022) and RMT Bulatov et al. (2022), we employ an uncased base BERT Devlin et al. (2018) as the feature extractor. Note that the batch size is set to 1 since we treat all instances in each single long document as a mini-batch to efficiently generate the feature vectors in parallel. Therefore, the actual batch size varies depending on the number of instances in the long document. We construct LaFiAttention layer with eight attention heads. With these settings, 100% of large documents from all six benchmark datasets can be fully processed using a single Tesla V-100 GPU with 32GB of memory on an NVIDIA DGX Station. As a result, the average inference time (0.026s) of each mini-batch is almost the same as BERT (0.022s). During training, the Adam optimizer Kingma and Ba (2014) is leveraged. As for the loss function, it varies depending on specific classification task. Following the baseline work Park et al. (2022), we use sigmoid and binary cross entropy for binary and multi-label classification, and softmax and cross entropy loss for multi-class classification. For Hyperpartisan Kiesel et al. (2019), Book Summary Bamman and Smith (2013) and EURLEX-57K Chalkidis et al. (2019), the learning rate 5e-6 is adopted. We find that the learning rate 5e-7 is more suitable for 20NewsGroups Lang (1995). We fine-tune the model for 20, 40, 60, and 100 epochs on Hyperpartisan, 20NewsGroups, EURLEX-57K, and Book Summary, respectively. We provide the experimental setup for code defect detection in the Appendix A.3.
Table 2: Performance metrics on only long documents in test set. The highest score in each column is bolded and underlined, while the second highest score is only bolded. The subsequent tables of this task are organized in a consistent manner.
@@ 5.1 Experimental Setup
Table 3: Performance metrics on full test set.
@@
Evaluation Metrics.
We evaluate the performance of LaFiCMIL using the same metrics as those employed in the baseline works Park et al. (2022); Hanif and Maffeis (2022). We report the accuracy (%) for both binary and multi-class classification. We use micro-F1 (%) for multi-label classification, which is based on summing up the individual true positives, false positives, and false negatives for each category. We report the detection accuracy (%) for code defect detection.
@@ 6 Experimental Results
In this section, we present and analyze the performance of the proposed LaFiCMIL in long document classification. We first discuss the overall performance, followed by an computational efficiency analysis and an ablation study on the core concepts of LaFiCMIL.
@@ 6.1 Overall Performance
Our experimental results reveal a phenomenon similar to  Park et al. (2022) in that no existing approach consistently outperforms the others across all benchmark datasets. However, as shown in Table 2, our LaFiCMIL establishes new state-of-the-art performance on all benchmark datasets when considering only long documents in the test set. Here, we define a long document as one containing at least two chunks (i.e., exceeding 512 BERT tokens). As shown in Table 3, we also achieve new state-of-the-art performance on five out of six benchmark datasets when considering the full data (i.e., a mix of long and short documents) in the test set. Particularly, we significantly improve the state-of-the-art score from 57.76% to 62.17% on the Paired Book Summary dataset, which contains the highest proportion of long documents (i.e., more than 75%). In contrast, we fail to achieve the best performance on 20NewsGroups, as the proportion of long documents in this dataset is very small (only 14.71%); thus, our improvement on long documents (as shown in Table 2) cannot dominate the overall performance on the entire dataset. This phenomenon is consistent with our motivation that the more large files (containing at least two chunks) present in the dataset, the more correlations LaFiCMIL can extract to boost classification performance.
Given that 100% of long documents from the six NLP benchmark datasets can be fully processed, we conduct an additional evaluation of LaFiCMIL’s ability to process extremely long sequences, based on the code defect detection dataset Devign. Our findings reveal that LaFiCMIL is capable of handling inputs of up to nearly 20 0002000020\,00020 000 tokens when utilizing CodeBERT Feng et al. (2020) and VulBERTa Hanif and Maffeis (2022) as feature extractors on a single GPU setup. This capability allows for 99.92% of the code files in the Devign dataset to be processed in their entirety. Concurrently, as demonstrated in Table 4, LaFiCMIL enhances the performance of both CodeBERT and VulBERTa, establishing a new state-of-the-art in accuracy over the evaluated baselines. Please find a detailed analysis on this task in the Appendix A.3.
Table 4: Accuracy (%) comparison of different models on the C programming language dataset for code defect detection. The highest accuracy score is bolded and underlined and the base model results are only bolded.
RoBERTa CodeBERT Code2vec PLBART VulBERTa CodeBERT+ LaFiCMIL VulBERTa+ LaFiCMIL
@@
Table 5: Runtime and memory requirements of each model, relative to BERT, based on experiments on the Hyperpartisan dataset. Training and inference time were measured and compared in seconds per epoch. GPU memory requirement is in GB.
@@
LaFiCMIL 1.86 1.18 32
@@ 6.3 Ablation Study
Without a feature extractor, any approach would be ineffective. Thus, when BERT is removed, the LaFiAttention layer must assume the role of feature extractor instead of c-MIL. This would result in the disappearance of the c-MIL mechanism, and the approach can now only take the first chunk as input, transforming it into a basic attention-based classifier. As might be expected, the absence of the BERT concept leads to the worst performance across all datasets among the three variants, as shown in the first row of Table 6. If excluding the LaFiAttention concept, c-MIL devolves into a standard MIL, for which we employ the widely accepted Attention-MIL Ilse et al. (2018). The results of this setting are presented in the second row of Table 6. Given that this variant can still process all chunks of a lengthy document, it performs best among all three variants on the four datasets with the largest number and longest length of documents. When the c-MIL concept is removed, the LaFiAttention layer will also be absent as it executes c-MIL which is no longer needed, leaving only BERT. Due to its restriction to process only the first chunk as input, this variant fails to achieve the best results on the four datasets with a preponderance of long documents. Finally, upon comparing the three variants with the full LaFiCMIL, shown in the fourth row of Table 6, it becomes evident that the exclusion of any concept significantly weakens performances across all datasets.
Furthermore, we observe some interesting results regarding chunk positional embedding. Basically, its effectiveness depends on different datasets. We perform an ablation study on chunk positional embedding, which is provided in the Appendix A.2.

### v4

@@ 4.2 LaFiCMIL
According to Theorem 1, we leverage Multi-layer Perceptron [40], BERT [13], LaFiAttention Layer and Layer Normalization [4] as sub-functions to approximate the c-MIL score function $S$ defined in Equation 10.
Given a set of bags $\{X_{1},...,X_{N}\}$, where each bag $X_{i}$ contains multiple instances $\{x_{i,1},...,x_{i,n}\}$, a bag label $Y_{i}$, and a randomly initialized category vector $x_{i,category}$, the goal is to learn the maps: $\mathbb{X}\rightarrow\mathbb{T}\rightarrow\gamma$, where $\mathbb{X}$ is the bag space, $\mathbb{T}$ is the transformer space and $\gamma$ is the label space. The map of $\mathbb{X}\rightarrow\mathbb{T}$ can be defined as:
$\small\begin{split}X_{i}^{0}=&[x_{i,category};f(x_{i,1});...;f(x_{i,n})]\\ &+E_{pos},\indent X_{i}^{0},E_{pos}\in\mathbb{R}^{(n+1)\times d}\end{split}$ (11)
$\small\begin{split}Q^{l}=X_{i}^{l-1}W_{Q},&\indent K^{l}=X_{i}^{l-1}W_{K},\indent V^{l}=X_{i}^{l-1}W_{V},\\ &\indent l=1,...,L\end{split}$ (12)
where function $f$ is approximated by a BERT model, $E_{pos}$ is the Positional Embedding, and $L$ is the number of Multi-head Self-Attention ($MSA$) block.
$\small\begin{split}head&=LaFiSA(Q^{l},K^{l},V^{l})\\ &=softmax(\frac{Q^{l}(\widetilde{K}^{l})^{T}}{\sqrt{d_{q}}})Z^{*l}softmax(\frac{\widetilde{Q}^{l}(K^{l})^{T}}{\sqrt{d_{q}}})V^{l},\\ \end{split}$ (13)
$\small MSA(Q^{l},K^{l},V^{l})=Concat(head_{1},...,head_{h})W_{O},$ (14)
$\small X_{i}^{l}=MSA(LN(X_{i}^{l-1}))+X_{i}^{l-1},\indent l=1,...,L$ (15)
where $W_{O}\in\mathbb{R}^{hd_{v}\times d}$, $head\in\mathbb{R}^{(n+1)\times d_{v}}$, $LaFiSA$ denotes the approximated Self-attention layer by Nyström method [5] according to Equation 8, $h$ is the number of head in each $MSA$ block, and Layer Normalization(LN) is applied before each $MSA$ block.
@@
From the above formulation, we can find that the most important part is to efficiently learn the map from bag space $\mathbb{X}$ to Transformer space $\mathbb{T}$. As illustrated in Figure 1, this map is approximated by a series of sub-functions which are approximated by various neural layers. The overall process is summarized as follows: given a large file, we use a BERT model to generate the representations of the divided chunks (i.e., instances in the concept of c-MIL). Then, we initialize a learnable category vector that follows a normal distribution and has the same shape as each instance. By considering the category vector as an additional instance, we learn the correlation between each instance using LaFiAttention layer. With the help of the attention mechanism, the category vector exchanges information with each chunk and extracts necessary features for large file classification. Finally, the category vector is fed into a fully connected layer to finalize the classification task.
5 Experimental Setup
@@
20NewsGroups multi-class 18846 10182 1132 7532 $20$ 368.83$\pm$783.84 $14.71$
Book Summary multi-label 12788 10230 1279 1279 $227$ 574.31$\pm$659.56 $38.46$
@@
EURLEX-57K multi-label 57000 45000 6000 6000 $4271$ 707.99$\pm$538.69 $51.3$
-Inverted multi-label 57000 45000 6000 6000 $4271$ 707.99$\pm$538.69 $51.3$
Devign binary 27318 21854 2732 2732 $2$ 615.46$\pm$41917.54 $39.76$
Table 2: Performance metrics on only long documents in test set. The highest score is bolded and underlined, while the second highest score is only bolded. The subsequent tables of this task are organized in a consistent manner.
@@
Datasets. To ensure a fair comparison with baselines, we adopt the same benchmark datasets utilized in the state-of-the-arts for long document classification [33]. We first evaluate LaFiCMIL on these six benchmark datasets: ① Hyperpartisan [22], a compact dataset encompassing 645 documents, designed for binary classification. ② 20NewsGroups [26], comprising 20 balanced categories and 11 8461184611\,84611 846 documents. ③ CMU Book Summary [6], tailored for multi-label classification, contains 12 7881278812\,78812 788 documents and 227 genre labels. ④ Paired Book Summary [33], formulated by combining pairs of documents from the CMU Book Summary dataset, features longer documents. ⑤ EURLEX-57K [11], a substantial multi-label classification dataset consisting of 57 0005700057\,00057 000 EU legal documents and $4271$ available labels. ⑥ Inverted EURLEX-57K [33], a modified version of EURLEX-57K dataset in which the order of sections is inverted, ensuring that core information appears towards the end of the document. To evaluate our method’s effectiveness with longer documents, we incorporated the Devign dataset [56] for code defect detection. It features documents that are substantially longer than those in the other six datasets, with lengths approaching or exceeding 20,000 tokens. Table 1 provides details of the datasets, covering metrics such as the average token count and its standard deviation, the maximum and minimum token counts, and the percentage of large documents, etc.
Implementation Details. We split a long text document into chunks (i.e., c-MIL instances), and follow the standard BERT input length (i.e., 512 tokens) for each chunk. To ensure a fair comparison, in line with the baselines, we employ BERT [13] for the first six datasets, and CodeBERT [16] and VulBERTa [17] for Devign, as the feature extractor. Since we treat all chunks in each long document as a mini-batch, the actual batch size varies depending on the number of chunks in the long document. We construct LaFiAttention layer with eight attention heads. Under these configurations, a single Tesla V-100 GPU with 32GB of memory on an NVIDIA DGX Station can fully process 100% of the large documents in the first six benchmark datasets and 99.92% of those from Devign. As a result, the average inference time (0.026s) of each mini-batch is almost the same as BERT (0.022s). During training, the Adam optimizer [23] is leveraged. The loss function varies depending on specific task. Following the baseline [33], we use sigmoid and binary cross entropy for binary and multi-label classification, and softmax and cross entropy loss for multi-class classification. For the Hyperpartisan, Book Summary, EURLEX-57K, and Devign datasets, a learning rate of 5e-6 is used, while 5e-7 is applied for the 20NewsGroups dataset. We fine-tune the model for 10, 20, 40, 60, and 100 epochs on Devign, Hyperpartisan, 20NewsGroups, EURLEX-57K, and Book Summary, respectively.
Table 3: Performance metrics on complete test set. The highest score in each column is bolded and underlined, while the second highest score is only bolded.
@@
Evaluation Setup. We evaluate the performance of LaFiCMIL using the same metrics as those employed in the baselines [33, 17]. We report the accuracy (%) for binary and multi-class classification. We use micro-F1 (%) for multi-label classification, and detection accuracy (%) for code defect detection. The results are the average of five independent runs, each with a different random seed. As we explained in the introduction, our work aims to address this challenge by focusing on using a more practical resource: a single GPU with 32 GB of memory, which is only a tenth of what RGPT demands (i.e., 320 GB). Therefore, given the distinct resource requirements and application contexts, comparing our approach with LLMs falls outside the scope of this study.
@@ 6 Experimental Results
In this section, we analyze the performance of LaFiCMIL in long document classification. We first discuss the overall performance, followed by an computational efficiency analysis and an ablation study on the core concepts of LaFiCMIL.
@@ 6.1 Overall Performance
Our experimental results reveal a phenomenon similar to  [33] in that no existing approach consistently outperforms the others across all benchmark datasets. However, as shown in Table 2, our LaFiCMIL establishes new state-of-the-art performance on all six NLP benchmark datasets when considering only long documents in the test set. Here, we define a long document as one containing over 512 BERT tokens. As shown in Table 3, we also achieve new state-of-the-art performance on five out of six NLP datasets when considering the full data (i.e., a mix of long and short documents) in the test set. Particularly, we significantly improve the state-of-the-art score from 57.76% to 62.17% on the Paired Book Summary dataset, which contains the highest proportion of long documents (i.e., more than 75%). In contrast, we fail to achieve the best performance on 20NewsGroups, as the proportion of long documents in this dataset is very small (only 14.71%); thus, our improvement on long documents (as shown in Table 2) cannot dominate the overall performance on the entire dataset. This phenomenon is consistent with our motivation that the more long documents present in the dataset, the more correlations LaFiCMIL can extract to boost classification performance.
Table 4: Accuracy (%) comparison of different models on Devign dataset for code defect detection. The highest accuracy score is bolded and underlined, and the base model results are only bolded.
RoBERTa CodeBERT Code2vec [3] PLBART [2] VulBERTa CodeBERT+ LaFiCMIL VulBERTa+ LaFiCMIL
@@
Table 5: Runtime and memory requirements of each model, relative to BERT, based on the Hyperpartisan dataset. Training and inference time were measured and compared in seconds per epoch. GPU memory requirement is in GB.
@@
LaFiCMIL 1.86 1.18 <32
Given that 100% of long documents from the six NLP datasets can be fully processed, we conduct an additional evaluation of LaFiCMIL’s ability to process extremely long sequences, based on the code defect detection dataset Devign. Our findings reveal that LaFiCMIL is capable of handling inputs of up to nearly 20 0002000020\,00020 000 tokens when utilizing CodeBERT [16] and VulBERTa [17] as feature extractors on a single GPU setup. This capability allows for 99.92% of the code files in the Devign dataset to be processed in their entirety. Concurrently, as shown in Table 4, LaFiCMIL enhances the performance of both CodeBERT and VulBERTa, establishing a new state-of-the-art in accuracy over the evaluated baselines. We find that code defect detection is a challenging task on which most existing state-of-the-art models struggle to achieve even a single percentage point improvement over previous models. Nonetheless, our LaFiCMIL helps CodeBERT gain 1.35 percentage points, representing a significant improvement.
6.2 Computational Efficiency Analysis
6.3 Ablation Study
@@
Without a feature extractor, any approach would be ineffective. Thus, when BERT is removed, the LaFiAttention layer must assume the role of feature extractor instead of c-MIL. This would result in the disappearance of the c-MIL mechanism, and the approach can now only take the first chunk as input, transforming it into a basic attention-based classifier. As might be expected, the absence of the BERT concept leads to the worst performance across all datasets among the three variants, as shown in the first row of Table 6. If excluding the LaFiAttention concept, c-MIL devolves into a standard MIL, for which we employ the widely accepted Attention-MIL [19]. The results of this setting are presented in the second row of Table 6. Given that this variant can still process all chunks of a lengthy document, it performs best among all three variants on the four datasets with the largest number and longest length of documents. When the c-MIL concept is removed, the LaFiAttention layer will also be absent as it executes c-MIL which is no longer needed, leaving only BERT. Due to its restriction to process only the first chunk as input, this variant fails to achieve the best results on the four datasets that are dominated by long documents. Finally, upon comparing the three variants with the full LaFiCMIL, shown in the fourth row of Table 6, it becomes evident that the exclusion of any concept significantly weakens performances across all datasets.